OpenStudy (anonymous):

Widget wonders produces widgets. They have found that the cost, c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it costs $30 to produce 2 widgets, $106 to produce 4 widgets, and $670 to produce 10 widgets. what's the total cost of producing 5 widgets?

OpenStudy (anonymous):

cost = aw^2 + bw + c solve for a, b, c: 30 = a(4) + b(2) + c 106 = a(16) + b(4) + c 670 = a(100) + b(10) + c cost of 5 = substitute w=5 into equation you found